The door code to get into a top secret laboratory is 6 digits. the first 3 digits of the code are all odd and the last 3 digits are all even. Digits can be used more than once. How many possible codes are there to gain access to this laboratory?
This is a Permutations and Combinations question.

Answer by jim_thompson5910(35256)

There are 10 single digit numbers (0,1,2,3,4,5,6,7,9,0)

There are 5 even single digit numbers (0,2,4,6,8)

There are 5 odd single digit numbers (1,3,5,7,9)


So there are 5*5*5 = 125 ways to construct the first part of the code and 5*5*5 = 125 ways to construct the second part.

In total, there are 125*125 = 15625 codes possible.